# Structuring Reasoning for Complex Rules Beyond Flat Representations

### v1

@@ Abstract
Zhihao Yang1,2thanks: Equal contribution, Ancheng Xu1,211footnotemark: 1, Jingpeng Li5, Liang Yan5, Jiehui Zhou5, Zhen Qin5, Hengyun Chang5, Ahmadreza Argha3, Hamid Alinejad-Rokny3, Minghuan Tan1, Yujun Cai4thanks: Corresponding authors, Min Yang122footnotemark: 2, 1Shenzhen Key Laboratory for High Performance Data Mining, Shenzhen Institutes of Advanced Technology, Chinese Academy of Sciences 2University of Chinese Academy of Sciences, 3School of Biomedical Engineering, UNSW Sydney 4University of Queensland, 5Alibaba Group zh.yang30@outlook.com, min.yang@siat.ac.cn
@@ 1 Introduction
Large language models (LLMs) (Zhao et al., 2025; Xiao et al., 2025) have exhibited remarkable capabilities across a wide range of natural language processing tasks (Vaswani et al., 2017; Brown et al., 2020), demonstrating significant potential in rule-intensive domains such as e-commerce content moderation (He et al., 2024), legal advisory services (Yue et al., 2024), and financial risk analysis (Wu et al., 2023b). These domains demand precise rule interpretation and rigorous implementation, as even minor errors in judgment can result in legal risks, financial losses, or serious damage to credibility. However, when processing complex instruction sets characterized by densely interdependent rules and nuanced semantic relationships (Ouyang et al., 2022), the performance of LLMs deteriorates significantly (Srivastava et al., 2022). For instance, when an advertisement must simultaneously convey the seller’s promotional message and strictly comply with advertising laws, LLMs often struggle to navigate the subtle interactions between overlapping rules, leading to inconsistent judgments (Xu et al., 2025). This limitation stems from two fundamental problems: (1) Flat rule processing, where models treat rule systems as unstructured text collections rather than hierarchically organized frameworks with explicit logical relationships; and (2) Reasoning divergence, where models are distracted by superficial features or misled by ambiguous cues (Ji et al., 2023), failing to identify the most relevant rules.
@@
Existing reasoning approaches have attempted to address complex reasoning challenges but fall short in rule-intensive scenarios. Chain-of-Thought (CoT) (Wei et al., 2022) and its variants fail to ensure that each step strictly adheres to logic and facts. Errors in any intermediate step can propagate and collapse the coherence of the entire reasoning chain (Lyu et al., 2023; Yang et al., 2024b). Tree-based methods like Tree-of-Thought (ToT) (Yao et al., 2024) explore multiple paths but lack systematic rule verification. These approaches share a common limitation: they lack mechanisms for hierarchical rule organization, structured evidence verification, and systematic conflict resolution when multiple rules apply simultaneously.
@@
This gap reveals a fundamental mismatch between current LLM reasoning approaches and the requirements of rule-intensive domains. Effective rule-based reasoning calls for a new framework that enables structured verification processes rather than relying on free-form exploration. To this end, we propose a novel method, the Dynamic Adjudication Template (DAT), illustrated in Figure 1. This method emulates the cognitive approach of human experts (Bilalić, 2017; Liao and Varshney, 2021). Rather than beginning with detailed computation, experts first construct a high-level problem-solving framework and then identify critical points for focused analysis. Unlike static prompting techniques that rely on fixed templates, DAT introduces a dynamic, structured reasoning paradigm. It actively guides the model through a structured three-step process of Qualitative Assessment, Evidence Gathering, and Adjudication. The process begins with a high-level assessment of the problem. The model then focuses on key, complex, and error-prone decision points most relevant to the question, enabling targeted analysis. These targeted insights are finally synthesized into a logically sound and comprehensive adjudication. In our experiments on a rule-intensive e-commerce dataset, DAT improved the overall accuracy of Qwen-2.5-7B (Qwen et al., 2025) from 34.11% to 62.49%. It also outperformed larger CoT-equipped models such as Qwen-Max and Deepseek-R1 (DeepSeek-AI et al., 2025) on several complex rule-based subtasks within the dataset.
@@
    We report substantial gains on rule-intensive benchmarks. Our method enables small, efficient models to outperform much larger LLMs using Chain-of-Thought reasoning on complex rule-based tasks. This paves the way for high-performance applications in low-resource settings. Preliminary results also suggest promising generalization to vision-language models (VLMs) (Liang et al., 2024; Li et al., 2023), indicating a fruitful direction for future research.
@@ 2.1 Complex Rule-based Reasoning Benchmark
LLMs and VLMs have been widely adopted in e-commerce applications such as content moderation, product recommendation, and search (Jiang et al., 2024; Palen-Michel et al., 2024). Nevertheless, enabling these models to make accurate judgments under complex and dynamic platform rules remains a significant challenge. To evaluate model capabilities for such tasks, both academia and industry have developed specialized benchmarks. Among these, EVADE stands out as the first multimodal benchmark for evasive content detection in Chinese e-commerce. Derived from authentic advertising regulations and annotated by domain experts, this dataset comprises 2,833 textual samples and 13,961 images across six product categories. EVADE evaluates model performance via two distinct tasks: Single-Violation Judgment and All-in-One Judgment. By embedding semantically overlapping rules within instructions, it requires reasoning over complex, long-context, rule-dense scenarios, closely aligning with our focus on complex rule interpretation. Research using EVADE has revealed limitations in current models’ capacity to comprehend and apply multi-layered rules, highlighting the necessity for our proposed approach.
@@ 2.2 Prompt-based Reasoning Methods with Large Language Models
To advance the reasoning capabilities of large language models, researchers have developed numerous prompt-based methodologies. Prominent among these are CoT prompting (Wei et al., 2022) and its derivatives, including Least-to-Most (Zhou et al., 2022a) and Decomposed Prompting (Khot et al., 2022), which decompose complex problems into sequential subtasks and have proven effective across diverse reasoning domains. However, the inherent structural fragility of linear reasoning approaches manifests as cascading error propagation during sequential processing stages, as demonstrated in Yang et al. (2024a). Emerging paradigms such as Tree-of-Thought and Graph-of-Thought (Besta et al., 2024) mitigate this issue through adaptive, non-linear heuristic search mechanisms. While these frameworks offer potential for improved reasoning fidelity, they introduce substantial computational costs and rely on handcrafted, domain-specific prompting strategies (Yang et al., 2024a). This rigid templating paradigm is particularly deficient for dynamic, evolving real-world environments. Within our problem domain, the governing rules exhibit both combinatorial complexity and temporal adaptation, rendering rule-category-specific manual prompt engineering impractical and misaligned with the dynamic nature of the challenge.
@@ 3 Methodology
In this section, we introduce our novel framework for rule-intensive applications. For a given query $Q$ and rules $\mathcal{R}$, the framework selects the optimal reasoning template $t^{*}$ from a predefined library $T_{K}$, which then guides the model in executing a structured reasoning process, as shown in Figure 2. To address the challenges of template generation, selection, and application, we propose a pipeline framework that consists of three primary components: 1) Three-Stage Structured Reasoning. This component serves as the core of our framework. It utilizes the selected template to execute a structured three-step reasoning process—Qualitative Analysis, Evidence Gathering, and Adjudication—to derive a logically grounded and rule-consistent result. This structured inference relies on high-quality, task-aligned templates, which are sourced and selected through the following two components. 2) Dynamic Template Library Construction. To support the reasoning process, this component systematically generates and validates a diverse repository of high-quality reasoning templates, which serve as the foundational knowledge base for the framework. 3) Adaptive Template Selection. Given a query, this component dynamically selects the optimal template from the library by evaluating both general performance and task-specific fitness. Detailed prompts for each component can be found in the appendix C.2.
@@ 3.2.1 Template Generation and Expansion
We introduce a systematic template generation pipeline using Gemini 2.5-Pro (Comanici et al., 2025). The process begins by creating 10 distinct seed templates ($T_{0}$) using only task context, without problem-specific details. These templates capture diverse analytical perspectives. To enhance coverage and robustness, we perform multi-stage augmentation through structured continuation (Zhou et al., 2022b; Madaan et al., 2023) and style transfer (Jin et al., 2022). Structured continuation begins by providing the model with the first several steps of each seed template—specifically, from step 1 to step n1n{-}1. The model is then instructed to complete the remaining steps in a consistent and logical manner, expanding $T_{0}$ into a structurally diverse set $T_{1}$. We then apply style transfer to $T_{1}$ to introduce linguistic and syntactic diversity. By not specifying the exact styles, we harness the model’s inherent understanding of language to produce a diverse and less biased set of augmented data. Specifically, the model is instructed to modify existing templates in $v$ distinct styles. This step modifies expressions while preserving the task relevance, yielding the final augmented set $T_{2}$ with improved stylistic diversity and task adaptability.
@@ 3.3 Adaptive Template Selection
In rule-intensive scenarios, ensuring absolute controllability and determinism is critical. Directly using large-model-generated reasoning templates risks violating domain-specific constraints (Zhang et al., 2023). To address this, we propose a Global-Local Template Selector that draws from a curated repository of pre-validated templates (Lewis et al., 2020), ensuring constraint compliance and process reliability. As shown on the left of Figure 2, the selector retrieves and validates templates to align with rule-intensive constraints. Our dual-evaluation framework assesses both the macro-level stability across tasks and adaptability to individual instances. We identify optimal templates through a weighted integration of global and local performance metrics (Hastie et al., 2009), with min-max normalization applied to both scores:
@@ 3.3.2 Local Selector $S_{2}$
Scoring with Local Selector $S_{2}$. Under the preference learning framework, we aim for the local selector $S_{2}$ to score all templates in the template library at once, enabling it to identify the most suitable template for a given query. To achieve this, we pair the current query with each template in the template library and pass them into $S_{2}$, which calculates the probability of selecting a specific template for the given query. Templates with higher probabilities are assigned higher scores. For a template $T_{i}$ consisting of $m$ tokens, we compute the average negative log-likelihood (NLL) Goodfellow et al. (2016) as follows:
@@ 4.1.2 Complex Instruction Dataset
Our experimental dataset is sourced from EVADE, containing 2,833 text and 13,961 image samples from real e-commerce platforms. The data focuses on six sensitive product categories: weight loss, diseases, height growth, body shaping, female care, and male care. Due to their direct impact on consumer welfare, these categories are governed by exceptionally detailed and strict policy rules derived from advertising laws. The logical complexity of these rules makes the dataset an ideal benchmark for evaluating a model’s reasoning abilities. For your reference, I have included the prompt for one of the tasks, diseases, in the appendix C.1.
@@ 4.4 Discussion
We extended DAT to high-capacity vision-language models (VLMs) and observed promising results. Without modifying the original pipeline, instructing VLMs to follow structured reasoning purely through textual templates proved effective on models with strong multimodal comprehension, such as Qwen-VL-Max (Bai et al., 2025), GPT-4.0 (OpenAI et al., 2024), and Gemini-2.5-Pro (Table 4). We focused on these large VLMs because they exhibit better understanding of complex textual instructions, which is essential for interpreting and executing structured templates. In contrast, smaller VLMs often show weaker performance in following detailed instructions (Liu et al., 2024; Wu et al., 2023a; Yang et al., 2023) and interpreting complex templates, likely due to their limited capacity. In our evaluation, this gap led to reduced effectiveness with DAT. We hypothesize that smaller VLMs lack sufficient understanding of and adherence to DAT, which limits their ability to deliver optimal performance. This highlights a potential avenue for improvement: by tailoring the design of templates for vision-language tasks, it may be possible to enhance the performance of smaller VLMs in future work.
Task Claude GPT-4o Qwen-VL-Max
@@
Table 4: Performance Comparison of VLM Models on Partial Accuracy. Abbreviations: Claude = Claude 3.0, GPT = GPT-4o, Qwen-VL = Qwen-VL-Max.
@@ Appendix A Detailed Performance of All Models with DAT
Appendix B Introduction of EVADE-Bench

### v2

@@ Abstract
Zhihao Yang1,2thanks: Equal contribution, Ancheng Xu1,211footnotemark: 1, Jingpeng Li5, Liang Yan5, Jiehui Zhou5, Zhen Qin5, Hengyu Chang5, Yukun Chen1,2, Longze Chen1,2, Ahmadreza Argha3, Hamid Alinejad-Rokny3, Minghuan Tan1, Yujun Cai4thanks: Corresponding authors, Min Yang122footnotemark: 2, 1Shenzhen Key Laboratory for High Performance Data Mining, Shenzhen Institutes of Advanced Technology, Chinese Academy of Sciences 2University of Chinese Academy of Sciences, 3School of Biomedical Engineering, UNSW Sydney 4University of Queensland, 5Alibaba Group zh.yang30@outlook.com, min.yang@siat.ac.cn
@@ 1 Introduction
Large language models (LLMs) Zhao et al. (2023); Xiao et al. (2025) have exhibited remarkable capabilities across a wide range of natural language processing tasks Vaswani et al. (2017); Brown et al. (2020), demonstrating significant potential in rule-intensive domains such as e-commerce content moderation Li et al. (2024b), legal advisory services Yang et al. (2024c), and financial risk analysis Wu et al. (2023b). These domains demand precise rule interpretation and rigorous implementation, as even minor errors in judgment can result in legal risks, financial losses, or serious damage to credibility. However, when processing complex instruction sets characterized by densely interdependent rules and nuanced semantic relationships Ouyang et al. (2022), the performance of LLMs deteriorates significantly Srivastava et al. (2023). For instance, when an advertisement must simultaneously convey the seller’s promotional message and strictly comply with advertising laws, LLMs often struggle to navigate the subtle interactions between overlapping rules, leading to inconsistent judgments Xu et al. (2025). This limitation stems from two fundamental problems: (1) Flat rule processing, where models treat rule systems as unstructured text collections rather than hierarchically organized frameworks with explicit logical relationships; and (2) Reasoning divergence, where models are distracted by superficial features or misled by ambiguous cues Ji et al. (2023), failing to identify the most relevant rules.
@@
Existing reasoning approaches have attempted to address complex reasoning challenges but fall short in rule-intensive scenarios. Chain-of-Thought (CoT) Wei et al. (2022) and its variants fail to ensure that each step strictly adheres to logic and facts. Errors in any intermediate step can propagate and collapse the coherence of the entire reasoning chain Lanham et al. (2023). Tree-based methods like Tree-of-Thought (ToT) Yao et al. (2023) explore multiple paths but lack systematic rule verification. These approaches share a common limitation: they lack mechanisms for hierarchical rule organization, structured evidence verification, and systematic conflict resolution when multiple rules apply simultaneously.
@@
This gap reveals a fundamental mismatch between current LLM reasoning approaches and the requirements of rule-intensive domains. Effective rule-based reasoning calls for a new framework that enables structured verification processes rather than relying on free-form exploration. To this end, we propose a novel method, the Dynamic Adjudication Template (DAT), illustrated in Figure 1. This method emulates the cognitive approach of human experts Liao and Varshney (2021). Rather than beginning with detailed computation, experts first construct a high-level problem-solving framework and then identify critical points for focused analysis. Unlike static prompting techniques that rely on fixed templates, DAT introduces a dynamic, structured reasoning paradigm. It actively guides the model through a structured three-step process of Qualitative Assessment, Evidence Gathering, and Adjudication. The process begins with a high-level assessment of the problem. The model then focuses on key, complex, and error-prone decision points most relevant to the question, enabling targeted analysis. These targeted insights are finally synthesized into a logically sound and comprehensive adjudication. In our experiments on a rule-intensive e-commerce dataset, DAT improved the overall accuracy of Qwen-2.5-7B Yang et al. (2024a) from 34.11% to 62.49%. It also outperformed larger CoT-equipped models such as Qwen-Max and Deepseek-R1 Guo et al. (2025) on several complex rule-based subtasks within the dataset.
@@
    We report substantial gains on rule-intensive benchmarks. Our method enables small, efficient models to outperform much larger LLMs using Chain-of-Thought reasoning on complex rule-based tasks. This paves the way for high-performance applications in low-resource settings. Preliminary results also suggest promising generalization to vision-language models (VLMs) Liang et al. (2024); Li et al. (2024a), indicating a fruitful direction for future research.
@@ 2.1 Complex Rule-based Reasoning Benchmark
LLMs and VLMs have been widely adopted in e-commerce applications such as content moderation, product recommendation, and search Jiang et al. (2025); Palen-Michel et al. (2024). Nevertheless, enabling these models to make accurate judgments under complex and dynamic platform rules remains a significant challenge. To evaluate model capabilities for such tasks, both academia and industry have developed specialized benchmarks. Among these, EVADE stands out as the first multimodal benchmark for evasive content detection in Chinese e-commerce. Derived from authentic advertising regulations and annotated by domain experts, this dataset comprises 2,833 textual samples and 13,961 images across six product categories. EVADE evaluates model performance via two distinct tasks: Single-Violation Judgment and All-in-One Judgment. By embedding semantically overlapping rules within instructions, it requires reasoning over complex, long-context, rule-dense scenarios, closely aligning with our focus on complex rule interpretation. Research using EVADE has revealed limitations in current models’ capacity to comprehend and apply multi-layered rules, highlighting the necessity for our proposed approach.
@@ 2.2 Prompt-based Reasoning Methods with Large Language Models
To advance the reasoning capabilities of large language models, researchers have developed numerous prompt-based methodologies. Prominent among these are CoT prompting Wei et al. (2022) and its derivatives, including Least-to-Most Zhou et al. (2022a) and Decomposed Prompting Khot et al. (2022), which decompose complex problems into sequential subtasks and have proven effective across diverse reasoning domains. However, the inherent structural fragility of linear reasoning approaches manifests as cascading error propagation during sequential processing stages, as demonstrated in Yang et al. (2024b). Emerging paradigms such as Tree-of-Thought and Graph-of-Thought Besta et al. (2024) mitigate this issue through adaptive, non-linear heuristic search mechanisms. While these frameworks offer potential for improved reasoning fidelity, they introduce substantial computational costs and rely on handcrafted, domain-specific prompting strategies Yang et al. (2024b). This rigid templating paradigm is particularly deficient for dynamic, evolving real-world environments. Within our problem domain, the governing rules exhibit both combinatorial complexity and temporal adaptation, rendering rule-category-specific manual prompt engineering impractical and misaligned with the dynamic nature of the challenge.
@@ 3 Methodology
In this section, we introduce our novel framework for rule-intensive applications. For a given query $Q$ and rules $\mathcal{R}$, the framework selects the optimal reasoning template $t^{*}$ from a predefined library $T_{K}$, which then guides the model in executing a structured reasoning process, as shown in Figure 2. To address the challenges of template generation, selection, and application, we propose a pipeline framework that consists of three primary components: 1) Three-Stage Structured Reasoning. This component serves as the core of our framework. It utilizes the selected template to execute a structured three-step reasoning process—Qualitative Analysis, Evidence Gathering, and Adjudication—to derive a logically grounded and rule-consistent result. This structured inference relies on high-quality, task-aligned templates, which are sourced and selected through the following two components. 2) Dynamic Template Library Construction. To support the reasoning process, this component systematically generates and validates a diverse repository of high-quality reasoning templates, which serve as the foundational knowledge base for the framework. 3) Adaptive Template Selection. Given a query, this component dynamically selects the optimal template from the library by evaluating both general performance and task-specific fitness. Detailed prompts for each component can be found in the appendix D.2.
@@ 3.2.1 Template Generation and Expansion
We introduce a systematic template generation pipeline using Gemini 2.5-Pro Comanici et al. (2025). The process begins by creating 10 distinct seed templates ($T_{0}$) using only task context, without problem-specific details. These templates capture diverse analytical perspectives. To enhance coverage and robustness, we perform multi-stage augmentation through structured continuation Zhou et al. (2022b); Madaan et al. (2023) and style transfer Jin et al. (2022). Structured continuation begins by providing the model with the first several steps of each seed template—specifically, from step 1 to step n1n{-}1. The model is then instructed to complete the remaining steps in a consistent and logical manner, expanding $T_{0}$ into a structurally diverse set $T_{1}$. We then apply style transfer to $T_{1}$ to introduce linguistic and syntactic diversity. By not specifying the exact styles, we harness the model’s inherent understanding of language to produce a diverse and less biased set of augmented data. Specifically, the model is instructed to modify existing templates in $v$ distinct styles. This step modifies expressions while preserving the task relevance, yielding the final augmented set $T_{2}$ with improved stylistic diversity and task adaptability.
@@ 3.3 Adaptive Template Selection
In rule-intensive scenarios, ensuring absolute controllability and determinism is critical. Directly using large-model-generated reasoning templates risks violating domain-specific constraints Huang et al. (2025). To address this, we propose a Global-Local Template Selector that draws from a curated repository of pre-validated templates Lewis et al. (2020), ensuring constraint compliance and process reliability. As shown on the left of Figure 2, the selector retrieves and validates templates to align with rule-intensive constraints. Our dual-evaluation framework assesses both the macro-level stability across tasks and adaptability to individual instances. We identify optimal templates through a weighted integration of global and local performance metrics John Lu (2010), with min-max normalization applied to both scores:
@@ 3.3.2 Local Selector $S_{2}$
Scoring with Local Selector $S_{2}$. Under the preference learning framework, we aim for the local selector $S_{2}$ to score all templates in the template library at once, enabling it to identify the most suitable template for a given query. To achieve this, we pair the current query with each template in the template library and pass them into $S_{2}$, which calculates the probability of selecting a specific template for the given query. Templates with higher probabilities are assigned higher scores. For a template $T_{i}$ consisting of $m$ tokens, we compute the average negative log-likelihood (NLL) LeCun et al. (2015) as follows:
@@ 4.1.2 Complex Instruction Dataset
Our experimental dataset is sourced from EVADE, containing 2,833 text and 13,961 image samples from real e-commerce platforms. The data focuses on six sensitive product categories: weight loss, diseases, height growth, body shaping, female care, and male care. Due to their direct impact on consumer welfare, these categories are governed by exceptionally detailed and strict policy rules derived from advertising laws. The logical complexity of these rules makes the dataset an ideal benchmark for evaluating a model’s reasoning abilities. For your reference, I have included the prompt for one of the tasks, diseases, in the appendix D.1.
@@ 4.4 Discussion
We extended DAT to high-capacity vision-language models (VLMs) and observed promising results. Without modifying the original pipeline, instructing VLMs to follow structured reasoning purely through textual templates proved effective on models with strong multimodal comprehension, such as Qwen-VL-Max Bai et al. (2025), GPT-4.0 Hurst et al. (2024), and Gemini-2.5-Pro (Table 4). We focused on these large VLMs because they exhibit better understanding of complex textual instructions, which is essential for interpreting and executing structured templates. In contrast, smaller VLMs often show weaker performance in following detailed instructions Liu et al. (2024); Wu et al. (2023a); Yang et al. (2023) and interpreting complex templates, likely due to their limited capacity. In our evaluation, this gap led to reduced effectiveness with DAT. We hypothesize that smaller VLMs lack sufficient understanding of and adherence to DAT, which limits their ability to deliver optimal performance. This highlights a potential avenue for improvement: by tailoring the design of templates for vision-language tasks, it may be possible to enhance the performance of smaller VLMs in future work.
Task Claude3.0 GPT-4o Qwen-VL-Max
@@
Table 4: Performance Comparison of VLM Models on Partial Accuracy.
@@ Appendix A Detailed Performance of All Models with DAT
Appendix B Formal Definition of Rule-Based Reasoning
We adopt the Four-Level Rule System from EVADE-Bench and cast detection as priority-constrained inference rather than flat pattern matching. Concretely, we model classification as satisfiability of layered constraints with override priorities.
B.1 Preliminaries and Notation
Let $I$ denote the input text and $\mathcal{C}=\{c_{1},\dots,c_{n}\}$ the set of semantic categories (e.g., disease, prohibited item). For each $c\in\mathcal{C}$, let $\mathcal{K}(c)$ be explicit keyword patterns and $\mathcal{S}(c)$ latent semantic patterns. Let $\mathcal{E}$ be auxiliary constraint patterns (e.g., treatment terms) and $\mathcal{M}$ exemption patterns (high-priority negatives). We use a unified matcher $\mathrm{match}(I,p)$ that returns True iff pattern $p$ (token, phrase, or semantic template) is satisfied by $I$.
B.2 Hierarchical Rule Definition
We define four rule levels with increasing inferential complexity.
First- and Second-Order (atomic/semantic).
$R_{1}(I,c)$ $:=\exists k\in\mathcal{K}(c):\mathrm{match}(I,k),$ (9)
$R_{2}(I,c)$ $:=\neg R_{1}(I,c)\ \wedge\ \exists s\in\mathcal{S}(c):\mathrm{match}(I,s).$ (10)
Third-Order (combinatorial).
Let $R_{12}(I,c):=R_{1}(I,c)\vee R_{2}(I,c)$. We require cross-category co-satisfaction (optionally with an auxiliary constraint):
$R_{3}(I):=\exists c_{i}\neq c_{j}\in\mathcal{C}:\ R_{12}(I,c_{i})\ \wedge\ R_{12}(I,c_{j})\ \wedge\ \big(\exists t\in\mathcal{E}:\mathrm{match}(I,t)\big).$ (11)
(If auxiliary co-occurrence is not required in a task, drop the last conjunct.)
Fourth-Order (exemption/override).
$R_{4}(I):=\exists m\in\mathcal{M}:\ \mathrm{match}(I,m).$ (12)
B.3 Priority-Aware Classification
We impose a strict priority $R_{4}\succ R_{3}\succ R_{1}\succ R_{2}$. Let $\mathcal{V}(I):=\{c\in\mathcal{C}:R_{12}(I,c)\}$. The classifier $F$ is
$F(I)=\begin{cases}\textsc{Safe}&\text{if }R_{4}(I),\\[2.0pt] \textsc{Violation}\big(\mathcal{V}(I)\big)&\text{else if }R_{3}(I),\\[2.0pt] \textsc{Violation}(c)&\text{else if }\exists c\in\mathcal{C}:\ R_{1}(I,c),\\[2.0pt] \textsc{Violation}(c)&\text{else if }\exists c\in\mathcal{C}:\ R_{2}(I,c),\\[2.0pt] \textsc{Safe}&\text{otherwise}.\end{cases}$ (13)
When multiple $c$ satisfy a clause, we report $\textsc{Violation}(\mathcal{V}(I))$ or resolve ties by a task-specific priority over categories.
Appendix C Introduction of EVADE-Bench